40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Jul 2 00:25:14 2013, gautam, Update, Green Locking, Universal PDH box tuning circuit.pdfmodel.pdf
    Reply  Wed Jul 10 01:27:44 2013, gautam, Update, Green Locking, Y-end Green PDH open-loop transfer function Green_PDH_measurement.pdfY-end_Green_PDH.pdf
Message ID: 8789     Entry time: Tue Jul 2 00:25:14 2013     Reply to this: 8817
Author: gautam 
Type: Update 
Category: Green Locking 
Subject: Universal PDH box tuning 

 [Koji, Annalisa, Gautam]

Annalisa noticed that over the weekend the Y-arm green PDH was locked to a sideband, despite not having changed anything on the PDH box (the sign switch was left as it was). On friday, we tried turning on and off some of the filters on the slow servo (C1ALS_Y_SLOW) which may have changed something but this warranted further investigation. We initially thought that the demodulation phase was not at the optimal value, and decided to try introducing some capacitances in the path from the function generator to the LO input on the universal PDH box. We modelled the circuit and determined that significant phase change was introduced by capacitances between 1nF and 100nF, so we picked out some capacitors (WIMA FKP) and set up a breadboard on which to try these out.

After some trial and error, Koji dropped by and felt that the loop was optimized for the old laser, the various loop parameters had not been tweaked since the new laser was installed. The following parameters had to be optimized for the new laser;

  • Servo gain
  • LO frequency
  • LO modulation depth
  • Demodulation phase

The setup was as follows: 

  • PDH box error signal to Oscilloscope CH1
  • Green PD output to Oscilloscope CH2
  • No capacitor between Function Generator and the PDH box
  • 0.1Hz triangle wave (30 counts amplitude) applied to ETMY via awggui (so as to sweep the cavity and see stronger, more regular TEM00 flashes)

The PDH error signal did not have very well-defined features, so Koji tweaked the LO frequency and the modulation depth till we got a reasonably well-defined PDH signal. Then we turned the excitation off and locked the cavity to green. The servo gain was then optimized by reducing oscillations in the error signal. Eventually, we settled on values for the Servo Gain, LO frequency and modulation depth such that the UGF was ~20kHz (determined by looking at the frequency of oscillation of the error signal on an Oscilloscope), and the PDH signal had well-defined features (while the cavity was unlocked). The current parameters are

  • LO frequency: 205.020 kHz
  • modulation depth: 0.032 Vpp

We then proceeded to find the optimal demodulation phase by simulating the circuit with various capacitances between the function generator and the PDH box (circuit diagram and plots attached). The simulation seemed to suggest that there was no need to introduce any additional capacitance in this path (introducing a 1nF capacitance added a phase-lag of ~90 degrees-this was confirmed as the error-signal amplitude decreased drastically when we hooked up a 1nF capacitor on our makeshift breadboard). In the current configuration, the LO is connected directly to the PDH box.

 

Misc Points:

  • The phase shifter in the PDH box is not connected: the IC in the box, JSPHS-26, is designed for operation in the range 18-26MHz. If necessary, an all-pass-filter could be introduced, with a tuneable rheostat to adjust the phase for our frequency range. Right now, turning the knob marked "LO phase angle" on the front panel doesn't do anything. The mixer on the PDH board is also not used for the same reason.
  • PSL shutter was closed sometime earlier this evening, because we suspected some IR light was reaching the Green PD on the y-endtable, and was influencing the error signal. Its back open now.
  • Useful information about the old y-end laser relevant to selecting the right LO frequency, modulation depth, and servo gain can be found here and in elog 2746 and subsequent replies, though the details of how the measurement were made aren't entirely clear. The idea is that the characteristics of the piezoelectric element in the laser has some characteristics which will determine the optimal LO frequency, modulation depth and servo gain.

 

To Do:

Now that we are reasonably confident that the loop parameters are optimal, we need to stabilise the C1ALS_Y_SLOW loop to stabilise the beat note itself. Appropriate filters need to be added to this servo.

 

Circuit Diagram: 50 ohm input impedance on the source, 50 ohm output impedance seen on the PDH box, capacitance varied between 1nF and 100nF in steps.

circuit.pdf

Plots for various capacitances: Gold-green trace (largest amplitude) direct from LO, other traces at input to PDH box.

model.pdf

ELOG V3.1.3-